Suppose that X is normally distributed with mean 95 and standard deviation 13.

A. What is the probability that X is greater than 113.72?
Probability =

B. What value of X does only the top 14% exceed?
X =
Answer #1

Solution:- Given that mean = 95, standard deviation = 13

A. P(X > 113.72) = P((X-mean)/sd > (113.72-95)/13)
= P(Z > 1.44)
= 1 − P(Z < 1.44)
= 1 − 0.9251
= 0.0749
Use the standard normal table to find the area below 1.44 :

| Z | 0.0 0.1 0.2 0.3 0.4 0.5 10.6 0.7 0.8 09 1.0 1.1 12 1.3 1.4 0.00 0.01 0.02 0.03 0.04 05 0.504 0.508 0512 0.516 0.5398 15

B. for Z = 1.08

X = mean + Z*Sd
= 95 + 1.08*13
= 109.04
Answer #2

where are you getting 1.08